OpenStudy (anonymous):

What is the arc length of a segment of a circle with a central angle of 165 degrees and a radius of 3 inches?

OpenStudy (anonymous):

S = (angle/360) * (2pi r)

OpenStudy (anonymous):

s = r θ where s = arc length r = radius = 3 inches θ = central angle in radians = ( Π /180)*165 s = 3(Π/180)*165 = 8.64 inches

OpenStudy (anonymous):

is it x=11pi/12 inches

OpenStudy (anonymous):

@victoriasunshine

OpenStudy (anonymous):

It would have to be 8.639 inches

OpenStudy (anonymous):

but those arnt my answers there fractions with pi in them

OpenStudy (anonymous):

@victoriasunshine

OpenStudy (anonymous):

oh what are your answers for the question

OpenStudy (anonymous):

or

OpenStudy (anonymous):

@victoriasunshine

OpenStudy (anonymous):

how many try's do u have on this

OpenStudy (anonymous):

only one ugh @victoriasunshine

OpenStudy (anonymous):

i think it would be the forth one \[s=11\pi/4 inches \]